A forklift is used to lift a large 704 kg crate a distance of 5.54 m in 6.12 s. What is the efficiency of the forklift if it takes a total of 8040 J/s of chemical energy in the form of burning fuel to complete this task?


1
Expert's answer
2021-04-29T07:27:55-0400

"P_1=\frac{mgh}{t}=\frac{704\cdot 9.8\cdot 5.54}{6.12}=6250~W,"

"\eta=\frac{P_1}{P_2}=\frac{6250}{8040}=0.78=78\%."